OpenStudy (anonymous):

Medal To reupholster a large sofa, an interior designer needs 12 1/2 yd of fabric that cost $5.43 per yard and 5 3/4 of fabric that cost $6,94 per yard. Find that total cost of the two fabrics.

OpenStudy (hockeychick23):

so first make 12 1/2 into a fraction and 5 3/4 into a fraction. So 12 1/2= 12.5 5 3/4= 5.75 (12.5*5.43)+(5.75*6.94)

OpenStudy (anonymous):

how would ou solve this

OpenStudy (hockeychick23):

ok so first we do 12.5*5.43 so what do you get when you multiply these?

OpenStudy (anonymous):

67.875

OpenStudy (hockeychick23):

Correct, now what is 5.75*6.94

OpenStudy (anonymous):

39.905

OpenStudy (hockeychick23):

good so what is 39.905+67.875?

OpenStudy (anonymous):

107.78

OpenStudy (hockeychick23):

correct, so your answer is $107.78

OpenStudy (anonymous):

thank you :)

OpenStudy (hockeychick23):

sure :)